OpenStudy (anonymous):

Please help! Solution set for: 4x-y=8 x+2y=11

OpenStudy (anonymous):

3,4

OpenStudy (anonymous):

Thanks guys!

OpenStudy (anonymous):

Multiply first equation by 2: 8x - 2y = 16 x + 2y = 11 Add them: 9x = 27 Find x and then y...

OpenStudy (callisto):

4x-y=8 => y = 4x -8 -(1) x+2y=11 -(2) Sub. (1) into (2) x + 2 (4x-8) = 11 9x - 16 = 11 x = 3 Sub. x=3 into (1) y = 4(3) - 8 = 4 So, (x,y) = (3,4) @KRAZZy Please also show your workings instead of just giving the final answer to the question.

OpenStudy (anonymous):

Thanks Callisto

OpenStudy (anonymous):

ok @Callisto i will next time